Required information The following information applies to the questions displayed below.) Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units the relevant range of production is 500 units to 1,500 units): Sales Variable expenses Contribution margin Fixed expenses 15,000 9,000 6, 000 3,120 2,880 Net operating income 2. What is the contribution margin ratio? in ratio

Required information The following information applies to the questions displayed below] Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units): Sales Variable expenses Contribution margin Fixed expenses Net operating income $ 15,000 9,000 6,000 3120 2,880 3. What is the variable expense ratio? Variableexpeo nse ratio

Required information The following information applies to the questions displayed below. Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units): Sales Variable expenses Contribution margin Fixed expenses Net operating income $ 15,000 9,000 6,000 3,120 $2,880 4. If sales increase to 1,001 units, what would be the increase in net operating income? (Round your answer to 2 decimal places.) in net

Required information [The following information applies to the questions displayed below] Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units): Sales Variable expenses Contribution margin Pixed expenses Net operating income s 15,000 9,000 6,000 3,120 2,880 5. If sales decline to 900 units, what would be the net operating income?

Required information [The following information applies to the questions displayed below. Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units): Sales Variable expenses Contribution margin Fixed expenses Net operating income 15,000 9,000 6,000 $ 2,880 6. If the selling price increases by $2 per unit and the sales volume decreases by 100 units, what would be the net operating income?
Answer #1

2)

Contribution Margin Ratio = Contribution Margin / Sales*100 = ($6000 / $15000)*100 = 40%

3)

Variable Expense Ratio = Variable Expense / Sales*100 = ($9000 / $15000)*100 = 60%

4.

Computation of net operating income if sale increase to 1001 units:

Current sales = 1000 units

Increased sales = 1001 units

Sale price per unit = $15

Variable cost per unit = $9

Particulars

Amount($)

Sales(1001*15)

15015

Variable expense(1001*9)

-9009

Contribution margin

6006

Fixed Expense

-3120

Net operating income

2886

Increase in Net operating income

$6

5.

Computation of net operating income if sale decline to 900 units:

Current sales = 1000 units

Decreased sales = 900 units

Sale price per unit = 15

Variable cost per unit = 9

Particulars

Amount($)

Sales(900*15)

13500

Variable expense(900*9)

-8100

Contribution margin

5400

Fixed Expense

-3120

Net operating income

2280

Net operating income

2280

6.

Computation of net operating income if selling price increase by $2 and sale unit’s decrease by 100 units:

Current sales = 1000 units

Decreased sales = 900 units

Sale price per unit = 15 +2 = 17

Variable cost per unit = 9

Particulars

Amount($)

Sales(900*17)

15300

Variable expense(900*9)

-8100

Contribution margin

7200

Fixed Expense

-3120

Net operating income

4080

Net operating income

$4080